OpenStudy (madeleine2):

Write the point-slope form of the line that passes through (5, 5) and is parallel to a line with a slope of 1/4. Include all of your work in your final answer.

OpenStudy (madeleine2):

@3mar

OpenStudy (madeleine2):

i tried to put it into a calculator but it said input is not valid

OpenStudy (3mar):

put the data correctly, please!

OpenStudy (madeleine2):

y−y1=m(x−x1) y-5=.25(x-5) i plugged it into the equation and thats how far i got

OpenStudy (madeleine2):

@3mar

OpenStudy (3mar):

yes I was waiting

OpenStudy (3mar):

can you use this and try? https://www.desmos.com/calculator/j9xfuztxcf

OpenStudy (madeleine2):

y=0.25x+b is the answer right

OpenStudy (3mar):

If you don;t mind, you can mul. by 4 both sides to eliminate 1/4

OpenStudy (3mar):

your answer is correct but b is?

OpenStudy (3mar):

@madeleine2

OpenStudy (mathmale):

Madeleine2: You must somehow calculate the value of b, the y-intercept. You already have enough info to do that: y−y1=m(x−x1) y-5=.25(x-5) i plugged it into the equation and thats how far i got